# Designs for Enabling Collaboration in Human-Machine Teaming via Interactive and Explainable Systems

### v1

@@ Abstract
Collaborative robots and machine learning-based virtual agents are increasingly entering the human workspace with the aim of increasing productivity and enhancing safety. Despite this, we show in a ubiquitous experimental domain, Overcooked-AI, that state-of-the-art techniques for human-machine teaming (HMT), which rely on imitation or reinforcement learning, are brittle and result in a machine agent that aims to decouple the machine and human’s actions to act independently rather than in a synergistic fashion. To remedy this deficiency, we develop HMT approaches that enable iterative, mixed-initiative team development allowing end-users to interactively reprogram interpretable AI teammates. Our 50-subject study provides several findings that we summarize into guidelines. While all approaches underperform a simple collaborative heuristic (a critical, negative result for learning-based methods), we find that white-box approaches supported by interactive modification can lead to significant team development, outperforming white-box approaches alone, and black-box approaches are easier to train and result in better HMT performance highlighting a tradeoff between explainability and interactivity versus ease-of-training. Together, these findings present three important directions: 1) Improving the ability to generate collaborative agents with white-box models, 2) Better learning methods to facilitate collaboration rather than individualized coordination, and 3) Mixed-initiative interfaces that enable users, who may vary in ability, to improve collaboration.
@@ 1 Introduction
Successful human-machine teaming (HMT) has long been sought after for its wide utility across potential applications, ranging from virtual agents such as “clippy” that provide on-demand support for improving documents to embodied robotic healthcare aides that can provide doctors with a helping hand Muoio (2019). While promising, achieving fluent HMT is challenging because interactions with humans can be incredibly complex due to the diversity across users Mataric (2018), human teammates benefit from explainable systems to support the development of mental models Paleja et al. (2021), and the lack of bidirectional communication (i.e., unclear how humans can “tell” a machine online to perform a desired behavior) Wright et al. (2022). In this paper, we transition from the conventional approach of crafting an HMT solution that aims for flawless out-of-the-box performance to a paradigm where end-users can actively interact with and program AI teammates, fostering a more dynamic and developmental interaction between humans and AI. Specifically, we explore enabling humans to perform user-specific modifications to a collaborative AI’s interpretable policy representation across repeated iterations of teaming episodes and provide a set of design guidelines to support team development in HMT drawn from a large-scale user study.
@@
Recently, data-driven techniques (e.g., imitation and reinforcement learning) have become popular in HMT, allowing for the generation of collaborative agent behavior without cumbersome manual programming Strouse et al. (2021); Carroll et al. (2019). footnotetext: DISTRIBUTION STATEMENT A. Approved for public release. Distribution is unlimited. This material is based upon work supported by the Under Secretary of Defense for Research and Engineering under Air Force Contract No. FA8702-15-D-0001. Any opinions, findings, conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the Under Secretary of Defense for Research and Engineering. Delivered to the U.S. Government with Unlimited Rights, as defined in DFARS Part 252.227-7013 or 7014 (Feb 2014). Notwithstanding any copyright notice, U.S. Government rights in this work are defined by DFARS 252.227-7013 or DFARS 252.227-7014 as detailed above. Use of this work other than as specifically authorized by the U.S. Government may violate any copyrights that exist in this work. However, these prior works utilize opaque, black-box models, limiting human’s ability to develop a shared mental model and maintain situational awareness Mathieu et al. (2000), crucial for high-performance teaming Salas et al. (1992). We posit that successful, real-world HMT is not feasible without the use of white-box methods, especially in safety-critical domains such as healthcare and manufacturing. Furthermore, collaborative interactions with machines have often lacked the ability to effectively learn with and adapt to human teammates in real-time Lake et al. (2016). In ad hoc human-human teams, effective teaming is often developed through an iterative process Tuckman (1965). Bi-directional communication is often a key component of this process, enabling the development of successful coordination strategies Salas et al. (2008). In our work, we build towards such a team development paradigm in HMT by 1) creating a pathway of bi-directional communication, utilizing interpretable policy representations as a mechanism to allow users to understand their machine teammates and allowing for explicit teammate policy modification through an interface (users can modify the machine’s tree-based policy via a GUI), and 2) allowing for the process of iterative mixed-initiative team development through repeated teaming episodes. We believe this paradigm is necessary because human-partnered systems need explainable components and adaptable systems. We provide the following contributions:
@@
  • We provide a case study regarding prior work in HMT Carroll et al. (2019); Strouse et al. (2021), finding that the generated machine behavior is unable to adapt to human-preferred strategies, and that high performance is typically driven by independent machine actions rather than collaboration, which can ultimately result in a higher team score.
@@
  • We create a novel InterpretableML architecture to support the creation of tree-based cooperative agent policies via reinforcement learning and a GUI to allow users to modify the AI’s behavior to their specifications. This capability is promising, enabling end-users to “go under-the-hood” of machine learning models and tune affordances or interactively and iteratively reprogram behavior.
@@ 2 Preliminaries
Here, we introduce prior work in HMT and Explainable AI, Overcooked-AI, Tuckman’s Model, and Markov Games.
@@
Human-Machine Teaming – The field of HMT is concerned with understanding, designing, and evaluating machines for use by or with humans Chen and Barnes (2014). A popular technique that has been used to produce collaborative AI agents is Reinforcement Learning (RL) Mnih et al. (2013), where researchers have concentrated efforts on reducing the dissimilarity between synthetic human training partners and testing with human end-users. Approaches that have achieved some success include utilizing human gameplay data to finetune simulated training partners to behave more human-like Carroll et al. (2019), which can be expensive, and training with a diverse-skilled population of synthetic partners to create an agent that can better generalize to non-expert end-users Strouse et al. (2021), which may bias the AI teammate to exhibit individualized strategies, as we display in Section 3. We note our work focuses on an interaction different from AI-assisted decision-making or decision support. Here, a human and an agent must collaborate across a series of timesteps, aiming to maximize a multifaceted joint objective function.
@@
Explainable AI – xAI is concerned with understanding and interpreting the behavior of AI systems Linardatos et al. (2021). In our work, we follow recent trends that show black-box methods paired with local explanations can be harmful Rudin (2018) and utilize interpretable, white-box tree-based models in a multi-agent sequential decision-making problem. These models have been shown to be beneficial in improving the user’s ability to simulate a decision-making model Tambwekar et al. (2021) and providing users with increased situational awareness over a teammate’s behavior in an HMT setting Paleja et al. (2021). While tree-based models can provide users insight into the model, the complexity of the tree-based model limits its utility Lipton (2018). While we note this as a potential weakness of utilizing tree-based models, effective state representations can provide a tradeoff between granular control and tree depth. Accordingly, we design our trees to reason over a state-space with high-level binary features and multi-step macro-actions, expanded on below. Furthermore, in our work, we explore a paradigm where a user can directly modify and visualize a tree-based AI teammate the user is interacting with after a teaming episode. Prior work in explainable debugging Kulesza et al. (2015) and robotics Paxton et al. (2017); Fogli et al. (2022) has explored similar paradigms, creating interactive systems that allow end-users to modify agent behavior to increase performance, but has not explored deploying tree-based models trained via RL in an HMT setting.
@@
Overcooked-AI – Overcooked-AI Carroll et al. (2019) is a testbed to evaluate human-AI interaction. Here, two agents are tasked with creating and delivering as many soups as possible within a given time. Achieving a high score requires agents to navigate a kitchen and repeatedly complete a set of sequential high-level actions, including collecting ingredients, placing ingredients in pots, cooking ingredients into a soup, collecting a dish, getting the soup, and delivering it. Both players receive the same score increase upon delivering the soup. We modify the original Overcooked-AI game to be a simultaneous-move game as opposed to the original formulation of allowing agents to perform actions asynchronously. This modification prevents the collaborative score metric from being dominated by super-human AI speed, causing the overall score to be more reliant upon effective collaboration and strategy. We provide details about the state and action space below and complete details in the appendix.
@@
Tuckman’s Model – Tuckman describes the different stages that a team goes through before reaching high performance, including “Forming”, “Storming”, “Norming” and “Performing,” often seeing a drop in performance as team members acclimate, followed by a rise as team members understand how to collaborate. Assuming that human-machine teams will follow similar stages to human-human teams, this paper looks into how we can support human-machine teams in reaching the Performing stage, where the team is achieving its full potential and exhibiting the highest level of cooperation. We provide a depiction of these stages as part of Figure 2.
Markov Game – We formulate our setting as a Markov Game Littman (1994), defined by a set of global states, $S_{1},S_{2}\in S$, a set of actions, $A_{1},A_{2}\in A$, transition function, $T:S\times A_{1}\times A_{2}\mapsto S$. and reward function $r_{i}:S\times A_{i}\mapsto\mathbb{R}$. Agent $i$ aims to maximize its discounted reward $R_{i}=\sum_{t=0}^{T}\gamma^{t}r_{i}^{t}$, where $\gamma\in[0,1]$ is a discount factor. For training, we utilize agent-agent collaborative training, which trains two separate agents jointly via single-agent PPO. We utilize PantheonRL Sarkar et al. (2022) for training our agents, incorporating our novel tree-based architecture (Section 4.1) into the codebase.
@@ 3 A Gap in Teaming Performance
In this section, we present two examples to display a gap in the quality of AIs in HMT. Specifically, we look at two recent approaches to produce collaborative AI agents Strouse et al. (2021); Carroll et al. (2019). We argue and display that the AIs trained via these approaches are rigid and exhibit individualized behaviors, missing out on collaborative teaming strategies that can ultimately result in higher team scores. We require AI agents that can effectively reach a consensus with humans on a teaming strategy that ultimately results in high performance. In cases where the human has a preferred strategy, the AI teammate should be able to support said strategy.
In Figure 1, we display the Coordination Ring scenario. A simple collaboration strategy (which we term “human-preferred”) in this domain is to utilize the counter to continuously pass objects, minimizing agent movement through efficient handoffs. To test a set of collaboration strategies, we utilize agents publicly available from Carroll et al. In Figure 1, we display a frame-by-frame of the human-preferred coordination strategy (Figure 1(a)) and AI-preferred coordination strategy (Figure 1(b)), which was a strategy where agents act individually to collect ingredients and place them in pots. The latter behavior was inferred through repeated play with the publicly-available AI. With the human-preferred strategy, the AI agent freezes for the majority of the game, creating an extremely frustrating and low-performing AI teammate. In this scenario, the human (green) picks up an ingredient and places it on the counter at the start of the game. The AI agent (blue), unfamiliar with this teaming strategy, freezes for approximately $80\%$ of the remaining episode before finally placing an onion in the pot. With the AI-preferred strategy, the human is able to successfully team with the AI, with each agent retrieving and placing ingredients while moving in a clockwise motion, but the strategy is not optimal or what the human prefers.
@@
We provide videos in the supplementary. As the AI produced in Carroll et al. is created via RL against human-like AI teammates, the generated behavior may not be ideal for the current teammate, especially if the current teammate’s preferred strategy was not present in the original training dataset used to create human-like AI training partners. This highlights a need for systems that take initial trained policies and adapt them to human-preferred behavior.
In a second example, we utilize the Optional Collaboration domain, displayed in Figure 4(b), which is also utilized in our human-subjects experiment. This domain was designed to incentivize collaboration, where creating mixed-ingredient dishes facilitated by agents passing ingredients across the central counter will result in a higher score per dish. Here, we program two intelligent deterministic heuristics: In the first, each agent acts completely individually, cooking single-ingredient dishes and serving. In the second, agents share ingredients, which costs additional timesteps, but are able to successfully cook mixed ingredient dishes. We find that the collaboration strategy achieves a $408$ team score, approximately $30\%$ more score compared to the individualized strategy of $306$. We find that trained policies under Ficticious Co-Play Strouse et al. (2021) exhibit similar team score to that of the individual coordination strategy and further, find that real human end-users collaborating with these agents are unable to far surpass the individual strategy score. As Strouse et al. trains an agent to work well with a population of agents, where approximately a third of the diverse-skilled population of agents used in training are completely random agents, we posit that the teammate agent must compensate and exhibit individualized behavior, limiting the algorithm’s ability to effectively learn effective team coordination strategies.
@@ 4.1 Interpretable Discrete Control Trees
Architecture Our IDCTs are based on differentiable decision trees (DDTs) Suárez and Lutsko (1999) – a neural network architecture that takes the topology of a decision tree (DT). DDTs contain decision nodes and leaf nodes; however, each decision node within the DDT utilizes a sigmoid activation function (i.e., a “soft” decision) instead of a Boolean decision (i.e., a “hard” decision). Each decision node, $i$, is represented by a sigmoid function, displayed as $y_{i}=(1+\exp(-\alpha(\vec{w}_{i}^{T}\vec{x}-b_{i})))^{-1}$. As this representation is difficult to interpret, Paleja et al. (2022) presented differentiable crispification, which recasts each decision node to split upon a single dimension of the input feature and translates the outcome of a decision node so that the outcome is a Boolean decision rather than a set of probabilities. This, in turn, allows for an interpretable forward propagation through the model that traces down a single branch of a tree as well as gradient flow afforded by the straight-through trick to update parameters of the neural tree model. We utilize this approach to learn interpretable tree-based teammate policies via RL.
@@
We initialize our IDCTs to be symmetric DTs with $N_{l}$ decision leaves and $N_{l}-1$ decision nodes. Each decision leaf is represented by a sparse categorical probability distribution over actions. At each timestep, a state variable is propagated through each decision node, split on a single decision rule, with the output being a Boolean causing the decision to proceed via the left or right branch until arrival at a leaf node. At each leaf node, we sample from the respective distribution to produce a macro-action (e.g., in Overcooked-AI, “get an onion” or “place ingredient on counter”). Further, we improve model predictability by applying an L1 norm loss over leaf node distributions to ensure sparsity, penalizing high entropy action distributions at a leafWhile utilizing deterministic AI policies may be easier to understand for users, we found these models could not converge to similar performance as the stochastic-leaf IDCT policies during training.. Importantly, the resultant representation after training is that of a simple decision tree with categorical probability distributions at each leaf node.
@@
Contextual Pruning As we focus on creating agents that must cooperate with and be interpreted by humans, we must limit the size of our tree-based models to a certain depth to promote user understanding. Analogous to the “lottery ticket hypothesis” in network training that supports the practicality of employing large models Frankle and Carbin (2018), a small tree with a limited number of sub-trees (lottery tickets) may not have the representational power to learn a high-performing policy. Thus, the ability to effectively train IDCTs is at odds with maintaining user readability and simulatability. Following recent literature in neural network pruning Liang et al. (2021), we design a post-hoc contextual pruning algorithm that allows us to simplify large IDCT models while precisely adhering to model behavior by accounting for:

### v2

@@ Abstract
Collaborative robots and machine learning-based virtual agents are increasingly entering the human workspace with the aim of increasing productivity and enhancing safety. Despite this, we show in a ubiquitous experimental domain, Overcooked-AI, that state-of-the-art techniques for human-machine teaming (HMT), which rely on imitation or reinforcement learning, are brittle and result in a machine agent that aims to decouple the machine and human’s actions to act independently rather than in a synergistic fashion. To remedy this deficiency, we develop HMT approaches that enable iterative, mixed-initiative team development allowing end-users to interactively reprogram interpretable AI teammates. Our 50-subject study provides several findings that we summarize into guidelines. While all approaches underperform a simple collaborative heuristic (a critical, negative result for learning-based methods), we find that white-box approaches supported by interactive modification can lead to significant team development, outperforming white-box approaches alone, and that black-box approaches are easier to train and result in better HMT performance, highlighting a tradeoff between explainability and interactivity versus ease-of-training. Together, these findings present three important future research directions: 1) Improving the ability to generate collaborative agents with white-box models, 2) Better learning methods to facilitate collaboration rather than individualized coordination, and 3) Mixed-initiative interfaces that enable users, who may vary in ability, to improve collaboration.
@@ 1 Introduction
SuccessfulDISTRIBUTION STATEMENT A. Approved for public release. Distribution is unlimited. This material is based upon work supported by the Under Secretary of Defense for Research and Engineering under Air Force Contract No. FA8702-15-D-0001. Any opinions, findings, conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the Under Secretary of Defense for Research and Engineering. Delivered to the U.S. Government with Unlimited Rights, as defined in DFARS Part 252.227-7013 or 7014 (Feb 2014). Notwithstanding any copyright notice, U.S. Government rights in this work are defined by DFARS 252.227-7013 or DFARS 252.227-7014 as detailed above. Use of this work other than as specifically authorized by the U.S. Government may violate any copyrights that exist in this work. human-machine teaming (HMT) has long been sought after for its wide utility across potential applications, ranging from virtual agents such as “clippy" that provide on-demand support for improving documents to embodied robotic healthcare aides that can provide doctors with a helping hand [29]. While promising, achieving fluent HMT is challenging because interactions with humans can be incredibly complex due to the diversity across users [26], human teammates benefit from explainable systems to support the development of mental models [31], and the lack of bidirectional communication (i.e., unclear how humans can “tell" a machine online to perform a desired behavior) [46]. In this paper, we transition from the conventional approach of crafting an HMT solution that aims for flawless out-of-the-box performance to a paradigm where end-users can actively interact with and program AI teammates, fostering a more dynamic and developmental interaction between humans and AI. Specifically, we explore enabling humans to perform user-specific modifications to a collaborative AI’s interpretable policy representation across repeated iterations of teaming episodes and provide a set of design guidelines to support team development in HMT drawn from a large-scale user study.
@@
Recently, data-driven techniques (e.g., imitation and reinforcement learning) have become popular in HMT, allowing for the generation of collaborative agent behavior without cumbersome manual programming [40, 5]. However, these prior works utilize opaque, black-box models, limiting human’s ability to develop a shared mental model and maintain situational awareness [27], crucial for high-performance teaming [37]. We posit that successful, real-world HMT is not feasible without the use of white-box methods, especially in safety-critical domains such as healthcare and manufacturing. Furthermore, collaborative interactions with machines have often lacked the ability to effectively learn with and adapt to human teammates in real-time [19]. In ad hoc human-human teams, effective teaming is often developed through an iterative process [43]. Bi-directional communication is often a key component of this process, enabling the development of successful coordination strategies [36]. In our work, we build towards such a team development paradigm in HMT by 1) creating a pathway of bi-directional communication, utilizing interpretable policy representations as a mechanism to allow users to understand their machine teammates and allowing for explicit teammate policy modification through an interface (users can modify the machine’s tree-based policy via a GUI), and 2) allowing for the process of iterative mixed-initiative team development through repeated teaming episodes. We believe this paradigm is necessary because human-partnered systems need explainable components and adaptable systems. We provide the following contributions:
@@
  • We provide a case study regarding prior work in HMT [5, 40], finding that the generated machine behavior is unable to adapt to human-preferred strategies, and that high performance is typically driven by independent machine actions rather than collaboration, which can ultimately result in a higher team score.
@@
  • We create a novel InterpretableML architecture to support the creation of tree-based cooperative agent policies via reinforcement learning and a GUI to allow users to modify the AI’s behavior to their specifications. This capability is promising, enabling end-users to “go under-the-hood" of machine learning models and tune affordances or interactively and iteratively reprogram behavior.
@@ 2 Preliminaries
Here, we introduce prior work in HMT and Explainable AI, our experimental domain, Overcooked-AI, a model of team development used to understand our findings, Tuckman’s Model, and the mathematical framework under which we generate agents, Markov Games.
@@
Human-Machine Teaming – The field of HMT is concerned with understanding, designing, and evaluating machines for use by or with humans [6, 44, 30]. A popular technique that has been used to produce collaborative AI agents is Reinforcement Learning (RL) [28], where researchers have concentrated efforts on reducing the dissimilarity between synthetic human training partners and testing with human end-users. Approaches that have achieved some success include utilizing human gameplay data to finetune simulated training partners to behave more human-like [5], which can be expensive, and training with a diverse-skilled population of synthetic partners to create an agent that can better generalize to non-expert end-users [40], which may bias the AI teammate to exhibit individualized strategies, as we display in Section 3. We note our work focuses on an interaction different from AI-assisted decision-making or decision support. Here, a human and an agent must collaborate across a series of timesteps, aiming to maximize a multifaceted joint objective function.
@@
Explainable AI – xAI is concerned with understanding and interpreting the behavior of AI systems [23]. In our work, we follow recent trends that show black-box methods paired with local explanations can be harmful [34] and utilize interpretable, white-box tree-based models in a multi-agent sequential decision-making problem. These models have been shown to be beneficial in improving the user’s ability to simulate a decision-making model [42] and providing users with increased situational awareness over a teammate’s behavior in an HMT setting [31]. While tree-based models can provide users insight into the model, the complexity of the tree-based model limits its utility [24]. While we note this as a potential weakness of utilizing tree-based models, effective state representations can provide a tradeoff between granular control and tree depth. Accordingly, we design our trees to reason over a state-space with high-level binary features and multi-step macro-actions, expanded on below. Furthermore, in our work, we explore a paradigm where a user can directly modify and visualize a tree-based AI teammate the user is interacting with after a teaming episode. Prior work in explainable debugging [18] and robotics [33, 9] has explored similar paradigms, creating interactive systems that allow end-users to modify agent behavior to increase performance, but has not explored deploying tree-based models trained via RL in a collaborative HMT setting. We provide a working definition of what we mean by “interpretable" within the Appendix Section G.
@@
Overcooked-AI – Overcooked-AI [5] is a testbed to evaluate human-AI interaction and has been used across HMT research concerned with collaboration [40], teammate identification [14], intention prediction [45], and behavior influence [16]. Here, two agents are tasked with creating and delivering as many soups as possible within a given time. Achieving a high score requires agents to navigate a kitchen and repeatedly complete a set of sequential high-level actions, including collecting ingredients, placing ingredients in pots, cooking ingredients into a soup, collecting a dish, getting the soup, and delivering it. Both players receive the same score increase upon delivering the soup. We modify the original Overcooked-AI game to be a simultaneous-move game as opposed to the original formulation of allowing agents to perform actions asynchronously. This modification prevents the collaborative score metric from being dominated by super-human AI speed, causing the overall score to be more reliant upon effective collaboration and strategy. We provide details about the state and action space below and complete details in the appendix.
@@
Tuckman’s Model – Tuckman [43] describes the different stages that a team goes through before reaching high performance, including “Forming", “Storming", “Norming" and “Performing," often seeing a drop in performance as team members acclimate, followed by a rise as team members understand how to collaborate. Assuming that human-machine teams will follow similar stages to human-human teams, this paper looks into how we can support human-machine teams in reaching the Performing stage, where the team is achieving its full potential and exhibiting the highest level of cooperation. We provide a depiction of these stages as part of Figure 2.
Markov Game – We formulate our setting as a Markov Game [25], defined by a set of global states, $S_{1},S_{2}\in S$, a set of actions, $A_{1},A_{2}\in A$, transition function, $T:S\times A_{1}\times A_{2}\mapsto S$. and reward function $r_{i}:S\times A_{i}\mapsto\mathbb{R}$. Agent $i$ aims to maximize its discounted reward $R_{i}=\sum_{t=0}^{T}\gamma^{t}r_{i}^{t}$, where $\gamma\in[0,1]$ is a discount factor. For training, we utilize agent-agent collaborative training, which trains two separate agents jointly via single-agent PPO. We utilize PantheonRL [38] for training our agents, incorporating our novel tree-based architecture (Section 4.1) into the codebase.
@@ 3 A Gap in Teaming Performance
In this section, we present two examples to display a gap in the quality of AIs in HMT. Specifically, we look at two recent approaches to produce collaborative AI agents [40, 5]. We argue and display that the AIs trained via these approaches are rigid and exhibit individualized behaviors, missing out on collaborative teaming strategies that can ultimately result in higher team scores. We require AI agents that can effectively reach a consensus with humans on a teaming strategy that ultimately results in high performance. In cases where the human has a preferred strategy, the AI teammate should be able to support said strategy.
In Figure 1, we display the Coordination Ring scenario. A simple collaboration strategy (which we term “human-preferred") in this domain is to utilize the counter to continuously pass objects, minimizing agent movement through efficient handoffs. To test a set of collaboration strategies, we utilize agents publicly available from Carroll et al. [5]. In Figure 1, we display a frame-by-frame of the human-preferred coordination strategy (Figure 1(a)) and AI-preferred coordination strategy (Figure 1(b)), which was a strategy where agents act individually to collect ingredients and place them in pots. The latter behavior was inferred through repeated play with the publicly-available AI. With the human-preferred strategy, the AI agent freezes for the majority of the game, creating an extremely frustrating and low-performing AI teammate. In this scenario, the human (green) picks up an ingredient and places it on the counter at the start of the game. The AI agent (blue), unfamiliar with this teaming strategy, freezes for approximately $80\%$ of the remaining episode before finally placing an onion in the pot. With the AI-preferred strategy, the human is able to successfully team with the AI, with each agent retrieving and placing ingredients while moving in a clockwise motion, but the strategy is not optimal or what the human prefers.
@@
As the AI produced by Carroll et al. [5] is created via RL teaming human-like AI teammates, the generated behavior may not be ideal for the current teammate, especially if the current teammate’s preferred strategy was not present in the original training dataset used to create human-like AI training partners. This highlights a need for systems that can explain strategies exhibited by trained agent policies and allow humans to adapt these pre-trained policies toward human-preferred behavior.
In a second example, we utilize the Optional Collaboration domain, displayed on the right-hand side of Figure 4(b), which is also utilized in our human-subjects experiment. This domain was designed to incentivize collaboration, where creating mixed-ingredient dishes facilitated by agents passing ingredients across the central counter will result in a higher score per dish. Here, we program two intelligent deterministic heuristics: In the first, each agent acts completely individually, cooking single-ingredient dishes and serving. In the second, agents share ingredients, which costs additional timesteps, but are able to successfully cook mixed ingredient dishes. We find that the collaboration strategy achieves a $408$ cumulative team score, approximately $30\%$ more score compared to the individualized strategy of $306$. However, we find that trained policies under Ficticious Co-Play [40] exhibit similar team score to that of the individual coordination strategy and further, find that real human end-users collaborating with these agents are unable to far surpass the individual strategy score. As Strouse et al. [40] trains an agent to work well with a population of agents, where approximately a third of the diverse-skilled population of agents used in training are completely random agents, we posit that the teammate agent must compensate and exhibit individualized behavior, limiting the algorithm’s ability to effectively learn effective team coordination strategies. In line with the first case study, the trained collaborative agent policies miss out on high-performance teaming behaviors, and thus, we need systems where humans can iteratively improve agent behavior online.
@@ 4.1 Interpretable Discrete Control Trees
Architecture Our IDCTs are based on differentiable decision trees (DDTs) [41] – a neural network architecture that takes the topology of a decision tree (DT). DDTs contain decision nodes and leaf nodes; however, each decision node within the DDT utilizes a sigmoid activation function (i.e., a “soft" decision) instead of a Boolean decision (i.e., a “hard" decision). Each decision node, $i$, is represented by a sigmoid function, displayed as $y_{i}=(1+\exp(-\alpha(\vec{w}_{i}^{T}\vec{x}-b_{i})))^{-1}$. As this representation is difficult to interpret, Paleja et al. [32] presented differentiable crispification, which recasts each decision node to split upon a single dimension of the input feature and translates the outcome of a decision node so that the outcome is a Boolean decision rather than a set of probabilities. This, in turn, allows for an interpretable forward propagation through the model that traces down a single branch of a tree as well as gradient flow afforded by the straight-through trick to update parameters of the neural tree model. We utilize this approach to learn interpretable tree-based teammate policies via reinforcement learning.
@@
We initialize our IDCTs to be symmetric DTs with $N_{l}$ decision leaves and $N_{l}-1$ decision nodes. Each decision leaf is represented by a sparse categorical probability distribution over actions. At each timestep, a state variable is propagated through each decision node, split on a single decision rule, with the output being a Boolean causing the decision to proceed via the left or right branch until arrival at a leaf node. At each leaf node, we sample from the respective distribution to produce a macro-action (e.g., in Overcooked-AI, “get an onion" or “place ingredient on counter"). Further, we improve model predictability by applying an L1 norm loss over leaf node distributions to ensure sparsity, penalizing high entropy action distributions at a leaf [^1] While utilizing deterministic AI policies may be easier to understand for users, we found these models could not converge to similar performance as the stochastic-leaf IDCT policies during training.. Importantly, the resultant representation after training is that of a simple decision tree with categorical probability distributions at each leaf node.
@@
Contextual Pruning As we focus on creating agents that must cooperate with and be interpreted by humans, we must limit the size of our tree-based models to a certain depth to promote user understanding. Analogous to the “lottery ticket hypothesis" in network training that supports the practicality of employing large models [11], a small tree with a limited number of sub-trees (lottery tickets) may not have the representational power to learn a high-performing policy. Thus, the ability to effectively train IDCTs is at odds with maintaining user readability and simulatability. Following work in neural network pruning [22], we design a post-hoc contextual pruning algorithm that allows us to simplify large IDCT models while precisely adhering to model behavior by accounting for: